OpenStudy (anonymous):

12(2x - 8) = 4(5 - 2x) + 3(7 - 3x) solve for x and leave in fraction form

OpenStudy (anonymous):

My answer was -55/23

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

Combine the terms on the LHS and the RHS. -96+24 x=41-17 x x (24 + 17) =41 + 96\[x=\frac{41+96}{24+17}=\frac{137}{41} \]